UT 29 May 2021 - up to 0.5 night MODS Checkout



Night Information

  • Sunset 19:25
  • Twilight Ends 20:58
  • Nightcenter 00:17
  • Handover from LBTO to LBTI 00:11

Night Plan

  • MODS Checkout - if for any reason MODS is NOT available tonight - proceeed with INAF Science programs (LUCI)
    • File location: /home/lbto/Service_Observing/2021A_D_E/May_Jun_Jul_2021/MODS_Checkout
    • Test running MODS 64bit GUIs on ROBS
    • AGWs guiding and collimation
    • AGW checks (if needed based on above)
    • ✅ MODS Astrometric Mask Rotator ZP check (10-20 min, about 10-min per one acquisition)
      • Stone L1 12 35 46 -01 01 01
      • MMS file mods.50.StoneI1.mms (/home/lbto/Service_Observing/2021A_D_E/May_Jun_Jul_2021/MODS_Checkout//RotZP/mods.50.StoneI1.mms)
    • ✅ DET XY offset test (15-20 detxy +/-Y offsets with Pointing Mask)
      • script uses dense portion of M13 field
    • Functional check - Spectrophotometric standards all modes (40-55 min)
      • Hz44 40-min total = 10 + 42-min or
        • ✅ Partially completed, executed in Grating modes only
      • GD153 52-min total = 10 + 42-min
    • Photometric Field M13 in dual, red-only and blue-imaging mode (20 min for 1 image per filter).
  • ✅ Reconfigure for LUCI - Critical AO OSCO check be executed before handover to LBTI
  • IF OSCO CANNOT BE EXCUTED:
    • Proceed with INAF Bino-LUCI SL Science Observations
    • Otherwise proceed with MODS Field Correction Check
      • MODS Field Correction check for IT 8315 -
        • run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection) do LDG and RDG in parallel (two IDL windows)
        • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
          • Stone_G is at RA=10:17,
          • Stone_H at RA=11:03
          • Stone_J at RA=13:32,
          • and other Stone / BS fields are available if earlier or later.

Night Log

Starting the night with binocular MODS

02:21 Steve is opening the enclosure.

Rotator Zero Point Checks: Statistics from 2020-2021 data:
inst mean median sig_n-1 n
MODS1 -0.17 -0.17 0.08 28
MODS2 -0.03 -0.04 0.07 30
02:50 Checking pointing and collimation on the Stone I field (airmass 1.2)

While collimating on-axis in twilight, it does look like the star may be a tad to the left of the pinhole on the MODS1 WFS image (UT 02:54). However, on the off-axis guide star used for the first field (StoneI1), the spots may be better centered (UT 03:05).

03:04 acqBinoMODS Stone_I1_acq.txt

Both the scripts hung with an error, after the slitmask it said it could not execute the command " " (a null command). This is odd. Trying to run the script piece-wise: acqMODS --mods1 -i Stone_I1_acq.txt and the same for MODS2, but it hung again on an empty/null command between "slitmask" and "go" commands which come from the single "slitGO" script command.
 [lbto@robs RotZP]$ acqMODS --mods2 -a Stone_I1_acq.txt

** Starting MODS2 acquisition script Stone_I1_acq.txt

** Executing the target acquisition image sequence without presetting
** the telescope or configuring the instrument.

** Executing MODS2 acquisition script Stone_I1_acq.txt

M2> slitmask in
DONE: SLITMASK SLITMASK=24 MASKPOS=IN MASKNAME='ID550828'
M2> 

03:19 Trying to run the script from obs3:

ssh -Y lbto@obs3
acqBinoMODS Stone_I1_acq.txt

works fine.

03:26 modsAlign is looking for the data in the local directory, needed to give the full path.
  • m1r: 20 and 21 --> offsetpointing -0.0951 -0.748 1.059 detxy rel
    • objects look aberrated & all a tad offset to the left
  • m2r: 20 and 21 --> offsetpointing -0.0299 3.883 -1.287 detxy rel
    • looks better

03:48
  • m1r: 23 and 24 --> offsetpointing -0.0938 -0.551 0.969 detxy rel
    • 26: centering of the stars in the boxes looks better this time, it was before 12 degon our first try. Collimation is better, but the stars are still elongated (at the center).
  • m2r: 23 and 24 --> offsetpointing -0.0253 3.935 -1.371 detxy rel
    • 26: pretty good as far as centering of stars in boxes.

03:57 acqBinoMODS hz44.acq
  • m1r: 26 offsetxy 0.376 11.141 rel --> 27 --> looks good
  • m2r: 26 offsetxy 5.940 7.602 rel --> 27 --> offsetxy -0.868 -0.044 rel --> 28

04:12 execBinoMODS hz44.obs
  • m1r: 28-30 (dual), 31-33 (red)
  • m1b: 14-16 (dual), 17-19 (blue)
  • m2r: 29-31 (dual), 32-34 (red)
  • m2b: 14-16 (dual), 17-19 (blue)
  • IT 8189 --- DX seeing is worse than SX. We're pointing to the NW and wind is from SW. Wind is <~8 m/s.
    • Steve closed the vent doors and that helped. The telescope is in equilibrium.
  • There is some haze on the horizon that we can see from the allsky, but the guide signals appear pretty steady
  • Observed only grating modes

04:51 execBinoMODS M13_pointing_abs.txt
  • FWHM is now worse 1.3 on SX and 0.8" on the DX side.
  • Pointing to the ENE and wind is from SW. Steve opened the vent door.
  • tweaking up starting position on mods2 :offsetxy 3.7 -2.5. One more offset to 0 +1
  • m1r: 35 - 53
  • m1b: 21 - 39
  • m2r: 38 - 57
  • m2b: 21 - 40
  • The first few images were taken with the FWHM ~1.3"/1" but now at 05:15 UT the FWHM on DX is bigger than on SX, but it is coming around again.
  • On MODS1, the WFS FWHM is 0.76" but the guider FWHM is 1-1.3 even, RMS WFE ~300nm. But there is a similar thing on MODS2: WFS FWHM ~ 0.8 but FWHM 1" * Do the MODS electronics produce dome seeing? Have we taken any near-IR images in the enclosure as is cools?

05:33 Finished with MODS checkout. Reconfiguring to bino-LUCI

The not-so-grand summary: MODS is functional, but there are some things that we need to look into:
  • Is the AGw hotspot correct for MODS1
  • The collimation on MODS1 was not as good as on MODS2. Images even near the center of the field looked elongated on the StoneI field. Some issues with dome seeing may have confused this.
  • The rotator zeropoints appeared to be ok - a bit off for MODS1 (0.08deg, just within errors) but were spot on for MODS2.
  • Took about 10 pairs of images dithering up/down with the Pointing Mask in place. I've downloaded them but will have to measure the direction of the offset with respect to the mask still.
  • Skipped photometric imaging fields. It was not truly photometric. It would be nice to get these at a later time since there is not much imaging data taken in photometric conditions (or it is time-consuming to identify the phot vs non-phot nights from the info we collect) and acq images are typically red-only.
  • MODS on robs was mostly successful. There were two issues which Yang tracked down on Saturday morning and I think more testing is warranted since after the initial hiccup with acqMODS, I ran the scripts from obs3.
    • The acqMODS script engine wrongly interpreted "slitGO" as "slitmask in" + "$go", not "slitmask in" + "go". "acqGO" was correctly interpreted and it seems this may have been an error in the commit to github, which would explain why it worked on obs3 (where it existed before being committed) and not on robs.
    • modsAlign expected the datafiles in the working directory. This is the default behavior if the environment variables MODS_IMGDIR and MODS_MMSDIR have not been set. On obsN, these are set in the ".cshrc" files.

5:41 Chevron pattern on OCAM. Cycled but persisted. wfsinit. When WFS came up the adsec went down. Xianyu took a core dump. Restarted adsec.

Fully cycling the wfs again, still have the chevron

wfsinit the housekeeper did not come up cleanly, no wfs arbitrator so started those up through system processes.

6:13 Finally up and running. Sending OSCO script.

Estimated wavefront magnitude in AOS 8.70 - some light cloud wisps. DIMM 1.05". NCPA on.
TN#062024, 36, 48
TN#062306,18,30

OSCO successfully completed. Results show successful communication for LUCI. We still have issues with WFS clobbering the adsec when brought up but that is not a new issue.

-- BarryRothberg - 28 May 2021
Topic revision: r13 - 30 May 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback